# Common origin of black holes in high mass X–ray binaries and in gravitational–wave sources

### v3

@@ Acknowledgements.
- Farmer et al. (2020) Farmer, R., Renzo, M., de Mink, S., Fishbach, M., & Justham, S. 2020, arXiv e-prints, arXiv:2006.06678
@@
- Fuller et al. (2019) Fuller, J., Piro, A. L., & Jermyn, A. S. 2019, MNRAS[arXiv:1902.08227]
@@
- Mroz et al. (2024) Mroz, P., Udalski, A., Szymanski, M. K., et al. 2024, arXiv e-prints, arXiv:2403.02386
@@ Appendix A Fitting Cyg X-1 spectra
We fit Cyg X-1 using two models: one considering a standard disc, and another where the disc is covered by some warm optically thick material (model a and b respectively). In both cases we model the spectra using xspec v.12.13.0 (Arnaud 1996).
@@
For the standard disc we use the xspec model kerrbb (Li et al. 2005). This considers multi-colour black-body emission originating from a disc structure, assuming Novikov & Thorne (1973) emissivity. All relativistic effects are taken into account, including ray-tracing from the disc to the observer. Additionally, this model always assumes that the disc extends to the ISCO.
@@
However, Cyg X-1 always shows a high energy tail that cannot be disc emission. This is generally understood as originating from a hot optically thin inner corona, which Compton up-scatters incident disc photons to higher energies. To model this we convolve our disc emission, kerrbb, with simpl (Steiner et al. 2009). This takes an input seed spectrum (in this case kerrbb) and scatters a fraction of the seed photons into a power-law component (hence emulating a Comptonized spectrum). In xspec syntax, the model is simpl*kerrbb.
@@
For model b we also have the addition of a warm optically thick skin covering the disc. This will also Compton up-scatter the disc photon to higher energies, but not sufficiently so to give a high energy tail. Instead this will give a spectrum that looks similar to a disc, but with the peak shifted to slightly higher energies and a shallower drop-off than the usual Wien tail. We model this using the convolution model thcomp (Zdziarski et al. 1996, 2020). This takes an input seed spectrum and Compton scatters to higher energies, similar to simpl, however unlike simpl thcomp takes the electron temperature of the Comptonizing plasma as an input argument, setting the high energy turn over. Additionally, for this optically thick skin we assume that all input seed photons are Compton scattered, unlike the case for the high-energy tail where only a fraction are scattered. In xspec syntax the model is now csimpl*thcomp*kerrbb.
@@
The line of sight absorption to Cyg X-1 is complex and variable. Hence we need to include an absorption component in both our models, for which we use tbabs. Due to the variable nature of the absorber in Cyg X-1 we leave the column-density in tbabs as a free parameter throughout. We note that in Fig. 3 we show the spectrum corrected for absorption in colour, and the original uncorrected spectrum in grey.
@@
As the data originate from two different instruments, XIS (soft-xrays) and HXD (hard-xrays), though both onboard Suzaku, we incldue a cross-calibration constant in all our models. Throughout this is fixed to 1.15, i.e increasing the model flux in the hard band by a factor 1.15. Finally, the total xspec models are: const*tbabs*csimpl*kerrbb for model a, and const*tbabs*csimpl*thcomp*kerrbb for model b.
Appendix B Re-Visiting M33-X7
@@
Here we re-visit the black hole spin estimate of M33-X7, performed by Liu et al. (2008), using the same methodology as for Cyg X-1. We extract the four Chandra ACIS spectra, referred to as the ’gold’ spectra in Liu et al. (2008). Initial inspection of the data show strongly disc dominated spectra, with no clear non-thermal high energy tail within the observed energy range. Hence we use a slightly simpler model than with Cyg X-1, modelling the continuum as either a pure disc (using kerrbb, Li et al. (2005)) or a disc with a warm Comptonising medium above it (using thcomp*kerrbb, Zdziarski et al. 1996, 2020). We also include the effects of absorption along our line of sight, which will affect the low energy emission, using tbabs. Hence the total xspec models are: tbabs*kerrbb for the pure disc case, and tbabs*thcomp*kerrbb for the warm Comptonised case.
@@
Fig. 4 shows the four ACIS spectra along with their respective model fits, for the pure disc (left) and warm Comptonised medium (right). The data show a change in normalisation between some of the spectra, corresponding to a change in the mass-accretion rate (see e.g the difference between the green/black data/model and the magenta/orange data/model in Fig. 4). Hence, during the spectral fitting we allow the mass-accretion rate to vary between the datasets, but then tie all other parameters as these should not vary between observations. The spectra were grouped such that each energy bin contained a minimum of 20 counts, allowing for the use of statistics.
Fig. 5 shows the change in as we step the fit through the possible values for black hole spin, with a standard disc on the left and a disc covered by a warm Comptonising material on the right. The standard disc clearly gives a strongly constrained value for the black hole spin, and in fact for the old mass estimate we recover the value derived in Liu et al. (2008) of . However, it is clear that a change in mass will affect this result, with the new mass estimate reducing the black hole spin to . However, the accretion flow is not necessarily characterised by a standard disc. Instead it is likely that it could be covered in a warm Comptonising material, as discussed previously for Cyg X-1 and in LMC X-1 (Zdziarski et al. 2024a, b). Unlike the case of a pure disc, this gives completely unconstrained black hole spin, with acceptable values ranging from -1 to at confidence. As with both Cyg X-1 and LMC X-1, this demonstrates the strongly model dependent nature of black hole spin estimates.

### v4

@@ Acknowledgements.
- Farmer et al. (2020) Farmer, R., Renzo, M., de Mink, S. E., Fishbach, M., & Justham, S. 2020, ApJ, 902, L36
@@
- Fuller et al. (2019) Fuller, J., Piro, A. L., & Jermyn, A. S. 2019, MNRAS, 485, 3661
@@
- Mróz et al. (2024) Mróz, P., Udalski, A., Szymański, M. K., et al. 2024, Nature, 632, 749
@@ Appendix A Fitting Cyg X-1 spectra
We fit Cyg X-1 using two models: 1) considering a standard disc and 2) where the disc is covered by some warm optically thick material (models a and b, respectively). In both cases we model the spectra using xspec v.12.13.0 (Arnaud 1996).
@@
For the standard disc we use the xspec model kerrbb (Li et al. 2005). This considers multi-colour blackbody emission originating from a disc structure, assuming Novikov & Thorne (1973) emissivity. All relativistic effects are taken into account, including ray-tracing from the disc to the observer. Additionally, this model always assumes that the disc extends to the ISCO.
@@
However, Cyg X-1 always shows a high energy tail that cannot be disc emission. This is generally understood as originating from a hot optically thin inner corona, which Compton up-scatters incident disc photons to higher energies. To model this we convolve our disc emission, kerrbb, with simpl (Steiner et al. 2009). This takes an input seed spectrum (in this case kerrbb) and scatters a fraction of the seed photons into a power-law component (hence emulating a Comptonised spectrum). In xspec syntax, the model is simpl*kerrbb.
@@
For model b we also have the addition of a warm optically thick skin covering the disc. This will also Compton up-scatter the disc photon to higher energies, but not sufficiently so to give a high energy tail. Instead this will give a spectrum that looks similar to a disc, but with the peak shifted to slightly higher energies and a shallower drop-off than the usual Wien tail. We model this using the convolution model thcomp (Zdziarski et al. 1996, 2020). This takes an input seed spectrum and Compton scatters to higher energies, similar to simpl, however unlike simpl thcomp takes the electron temperature of the Comptonising plasma as an input argument, setting the high energy turn over. Additionally, for this optically thick skin we assume that all input seed photons are Compton scattered, unlike the case for the high-energy tail where only a fraction are scattered. In xspec syntax the model is now csimpl*thcomp*kerrbb.
@@
The line of sight absorption to Cyg X-1 is complex and variable. Hence we need to include an absorption component in both our models, for which we use tbabs. Due to the variable nature of the absorber in Cyg X-1 we leave the column-density in tbabs as a free parameter throughout. We note that in Fig. 3 we show the spectrum corrected for absorption in colour and the original uncorrected spectrum in grey.
@@
As the data originate from two different instruments, XIS (soft-xrays) and HXD (hard-xrays), while both onboard Suzaku, we included a cross-calibration constant in all our models. Throughout, this was fixed to 1.15, that is: increasing the model flux in the hard band by a factor 1.15. Finally, the total xspec models are: const*tbabs*csimpl*kerrbb for model a, and const*tbabs*csimpl*thcomp*kerrbb for model b.
Appendix B Revisiting M33-X7
@@
Here we revisit the black hole spin estimate of M33-X7, performed by Liu et al. (2008), using the same methodology as for Cyg X-1. We extract the four Chandra ACIS spectra, referred to as the ’gold’ spectra in Liu et al. (2008). An initial inspection of the data show strongly disc dominated spectra, with no clear non-thermal high energy tail within the observed energy range. Hence, we used a slightly simpler model than with Cyg X-1, modelling the continuum as either a pure disc (using kerrbb, Li et al. (2005)) or a disc with a warm Comptonising medium above it (using thcomp*kerrbb, Zdziarski et al. 1996, 2020). We also include the effects of absorption along our line of sight, which will affect the low energy emission, using tbabs. Hence the total xspec models are: tbabs*kerrbb for the pure disc case, and tbabs*thcomp*kerrbb for the warm Comptonised case.
@@
Figure 4 shows the four ACIS spectra along with their respective model fits, for the pure disc (left) and warm Comptonised medium (right). The data show a change in normalisation between some of the spectra, corresponding to a change in the mass-accretion rate (see e.g the difference between the green-black data model and the magenta-orange data model in Fig. 4). Hence, during the spectral fitting we allow the mass-accretion rate to vary between the datasets, but then tie all other parameters as these should not vary between observations. The spectra were grouped such that each energy bin contained a minimum of 20 counts, allowing for the use of statistics.
Figure 5 shows the change in as we step the fit through the possible values for black hole spin, with a standard disc on the left and a disc covered by a warm Comptonising material on the right. The standard disc clearly gives a strongly constrained value for the black hole spin, and in fact for the old mass estimate we recover the value derived in Liu et al. (2008) of . However, it is clear that a change in mass will affect this result, with the new mass estimate reducing the black hole spin to . However, the accretion flow is not necessarily characterised by a standard disc. Instead it is likely that it could be covered in a warm Comptonising material, as discussed previously for Cyg X-1 and in LMC X-1 (Zdziarski et al. 2024a, b). Unlike the case of a pure disc, this gives completely unconstrained black hole spin, with acceptable values ranging from -1 to at confidence. As with both Cyg X-1 and LMC X-1, this demonstrates the strongly model dependent nature of black hole spin estimates.